# Is fixed-node diffusion quantum Monte Carlo reproducible?

### v2

@@ Abstract
Fixed-node diffusion quantum Monte Carlo (FN-DMC) is a widely-trusted many-body method for solving the Schrödinger equation, known for its reliable predictions of material and molecular properties. Furthermore, its excellent scalability with system complexity and near-perfect utilization of computational power makes FN-DMC ideally positioned to leverage new advances in computing to address increasingly complex scientific problems. Even though the method is widely used as a computational gold standard, reproducibility across the numerous FN-DMC code implementations has yet to be demonstrated. This difficulty stems from the diverse array of DMC algorithms and trial wave functions, compounded by the method’s inherent stochastic nature. This study represents a community-wide effort to address the titular question, affirming that: Yes, FN-DMC is reproducible (when handled with care). Using the water-methane dimer as the canonical test case, we compare results from eleven different FN-DMC codes and show that the approximations to treat the non-locality of pseudopotentials are the primary source of the discrepancies between them. In particular, we demonstrate that, for the same choice of determinantal component in the trial wave function, reliable and reproducible predictions can be achieved by employing the T-move (TM), the determinant locality approximation (DLA), or the determinant T-move (DTM) schemes, while the older locality approximation (LA) leads to considerable variability in results. This work lays the foundation to establish accurate and reproducible FN-DMC estimates for all future studies across applications in materials science, physics, chemistry, and biology.
@@ 1 Introduction
Fixed-node diffusion Monte Carlo[27, 28, 29] (FN-DMC) is an accurate state-of-the-art computational approach for solving the Schrödinger equation for a variety of systems, including molecules, solids, and surfaces. This method obtains the electronic ground-state by performing an imaginary-time evolution from a starting trial wave function . Within the Born-Oppenheimer approximation, consists of the real space positions of all the electrons. Typically, is the product of an antisymmetric function (e.g., a Slater determinant or a sum of Slater determinants[30]) and a symmetric, positive function, called the Jastrow factor[31]. The Jastrow factor is explicitly dependent on electron-electron and electron-nucleus distances, and able to directly capture a significant fraction of the electronic correlation.
@@
There are numerous QMC codes currently used for research, many of which have been under development for over a decade. Each makes somewhat different algorithmic and implementation choices, such as use of different Jastrow factors and methods for evaluating single particle orbitals. As detailed below, we compare eleven of them in this study. This diversity raises open questions on the reproducibility of FN-DMC. If FN-DMC is to be widely accepted as a highly accurate reference method, it is important that consistent results can be obtained from these different FN-DMC codes. With this goal in mind, the present letter represents a collaborative effort among the users and developers of eleven distinct FN-DMC codes, to rigorously assess the reproducibility of FN-DMC.
@@
A key obstacle to the reproducibility of FN-DMC comes from the use of non-local pseudopotentials (NLPPs), which increase the efficiency of the method for systems with heavy atoms. [95, 96, 97] NLPPs allow one to solve the Schrödinger equation solely for the valence electrons, by substituting the full local nuclear potential with a smooth non-local potential near the nuclei. In general, NLPPs hinder reproducibility in electronic structure methods, as NLPPs constructed in different ways can lead to somewhat different predictions. NLPPs are a potential source of non-reproducibility in FN-DMC even when the same NLPPs are used, because non-local pseudopotential operators create an additional sign problem in the projector beyond the one that is always present for fermionic calculations. To avoid this sign problem, these operators must be “localized”,[98] or at least partially localized,[99] on a wave-function. A natural choice is to localize them on the trial wave-function , introducing a dependence on both the determinantal and the Jastrow components of the wave function. Since the Jastrow factor is different in the different codes and its parameters are stochastically optimized, yielding possible noise and reproducibility issues, some authors choose to localize only on the determinantal component. [100, 96, 101, 102, 103] This removes the dependence on the Jastrow factor at the cost of losing the desirable property that the treatment of the pseudopotential is exact in the limit of exact . To summarize, there are currently four localization schemes: the locality approximation (LA)[104, 98], the T-move (TM) approximation[99, 105, 106], the determinant locality approximation (DLA)[100, 96, 101, 102, 103], and the determinant T-move (DTM) approximation[103]. These four schemes (LA, TM, DLA and DTM) result in somewhat different projected wave-functions and therefore different total energies of physical systems.
In this work, we systematically compare the four localization algorithms across eleven FN-DMC codes (named alphabetically): Amolqc, CASINO[34], CHAMP-EU[107], CHAMP-US[108], CMQMC, PyQMC[109], QMC=Chem[110, 111], QMCPACK[36, 112], QMeCha[113], QWalk[114], and TurboRVB[35, 115]. Different forms of Jastrow factor are necessarily tested as part of this evaluation. We specifically consider the case of the interaction energy between methane and water molecules. The methane-water dimer is characterized by weak binding (), which makes it a highly stringent target and thus a great test for reproducibility. In addition, it is a prototype of more complex systems such as methane clathrates, important for gas storage and transportation. We show that consensus across all eleven codes can be made when utilizing the TM, DLA and DTM approximations, particularly following careful control of the discretized time step.
@@ 2 Results and Discussion
is defined as the difference between the energy of the complex, , minus the sum of the energies of the isolated water and methane monomers (see the Methods section for details on the geometries and the DMC simulation set-up). All the interaction energies are extrapolated to the zero time-step limit according to the procedure described in the supporting information (SI) and in Ref. [116].
@@ 3 Methods
The interaction energy of the methane-water dimer is computed by subtracting the isolated molecule energies from the methane-water complex, as defined in Eq. 1. The geometry of the dimer (shown in the SI) was obtained from Ref. [122]. The geometries of the monomers are the same as in the dimer. In this study, in order to try to achieve consistent results, all eleven codes were required to use the same correlation consistent effective core potential (ccECPs)[123, 124] and the corresponding triple-zeta basis set (ccECP-ccpVTZ), as well as a Slater-Jastrow wave function with a single Slater determinant whose orbitals are obtained from DFT calculations using the Perdew-Zunger parametrization[125] of the local-density approximation. For the methane-water dimer, this was sufficient to obtain accurate results. Some of the codes exchanged wave function data via the TREXIO library. [126] Every code implements a slightly different parametrization of the Jastrow factor, but all codes include in the Jastrow factor an electron-electron (e-e), an electron-nucleus (e-n), and an electron-electron-nucleus (e-e-n) term. The variational parameters of the Jastrow factor have been optimized by minimizing either the variational energy or the variance, according to the recommended scheme within each code. The time steps employed in each simulation are in the range 0.001 to 0.1 a.u. The final estimates reported in Fig. 1 were extrapolated to the limit using the procedure described in the SI. Further details specific to each code, the schemes used to deal with the localization error, the time step extrapolation, and the tests on the size consistency error are reported in the SI.
@@ 4 Summary and Conclusions
In particular, we conducted a thorough analysis of two key obstacles to FN-DMC reproducibility, namely the use of NLPPs and finite time-step bias. We systematically compared four localization schemes, LA, TM, DLA, and DTM, for the interaction energy of the methane-water dimer and found that agreement across all eleven codes is achieved in the limit of zero time step when employing the TM, DLA, and DTM approximations. Notably, we achieve agreement within a standard deviation of meV on the interaction energy of the methane-water complex, approximately two hundred thousand times smaller than the total energy of the dimer. Larger discrepancies are observed with the LA scheme. Moreover, we show that even the total energies with the TM, DLA, and DTM schemes have a standard deviation among the codes which is smaller than meV. This agreement further reinforces the reproducibility of FN-DMC.
@@ Acknowledgements
M.B. acknowledges the computational resources from the HPC facilities of the University of Luxembourg [127] (see hpc.uni.lu). M.D. acknowledges financial support from the European Union under the LERCO project number CZ.10.03.01/00/22_003/0000003, via the Operational Programme Just Transition, and computational resources from the IT4Innovations National Supercomputing Center (e-INFRA CZ, ID: 90140). M.C. acknowledges access to French computational resources at the CEA-TGCC center through the GENCI allocation number A0150906493. J.T.K., P.R.C.K., Y.L. and L.M. were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. Part of the work of L.M. has been supported also by U.S. National Science Foundation grant DMR-2316007 and employed resources at NERSC at early stages for this project. M.C., E.S., R.S., and C.F. acknowledge partial support by the European Centre of Excellence in Exascale Computing TREX — Targeting Real Chemical Accuracy at the Exascale. This project has received funding in part from the European Union’s Horizon 2020 — Research and Innovation program — under grant agreement no. 952165. E.S., R.S., and C.F performed the calculations on the Dutch national supercomputer Snellius with the support of SURF Cooperative. K.N. acknowledges financial support from the JSPS Overseas Research Fellowships and from MEXT Leading Initiative for Excellent Young Researchers (Grant No. JPMXS0320220025) and computational resources from the Numerical Materials Simulator at National Institute for Materials Science (NIMS). L.K.W and W.W. were supported by U.S. National Science Foundation via Award No. 1931258. A.Z. and D.A. acknowledges support from the European Union under the Next generation EU (projects 20222FXZ33 and P2022MC742). Y.S.A., D.A. and A.Z. acknowledge support from Leverhulme grant no. RPG-2020-038. A.M. and B.X.S. acknowledge support from the European Union under the “n-AQUA” European Research Council project (Grant No. 101071937). The portion of the work done by T.A.A. and C.J.U. received initial support under AFOSR (Grant No. FA9550-18-1-0095) and was completed under the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration."
@@
This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Calculations were also performed using the Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk), provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/T022159/1 and EP/P020259/1), and DiRAC funding from the Science and Technology Facilities Council (www.dirac.ac.uk). This work also used the ARCHER UK National Supercomputing Service (https://www.archer2.ac.uk), the United Kingdom Car Parrinello (UKCP) consortium (EP/ F036884/1).

### v3

# Reproducibility of fixed-node diffusion Monte Carlo across diverse community codes: The case of water-methane dimer

@@ Abstract
Fixed-node diffusion quantum Monte Carlo (FN-DMC) is a widely-trusted many-body method for solving the Schrödinger equation, known for its reliable predictions of material and molecular properties. Furthermore, its excellent scalability with system complexity and near-perfect utilization of computational power makes FN-DMC ideally positioned to leverage new advances in computing to address increasingly complex scientific problems. Even though the method is widely used as a computational gold standard, reproducibility across the numerous FN-DMC code implementations has yet to be demonstrated. This difficulty stems from the diverse array of DMC algorithms and trial wave functions, compounded by the method’s inherent stochastic nature. This study represents a community-wide effort to assess the reproducibility of the method, affirming that: Yes, FN-DMC is reproducible (when handled with care). Using the water-methane dimer as the canonical test case, we compare results from eleven different FN-DMC codes and show that the approximations to treat the non-locality of pseudopotentials are the primary source of the discrepancies between them. In particular, we demonstrate that, for the same choice of determinantal component in the trial wave function, reliable and reproducible predictions can be achieved by employing the T-move (TM), the determinant locality approximation (DLA), or the determinant T-move (DTM) schemes, while the older locality approximation (LA) leads to considerable variability in results. These findings demonstrate that, with appropriate choices of algorithmic details, fixed-node DMC is reproducible across diverse community codes—highlighting the maturity and robustness of the method as a tool for open and reliable computational science.
@@ 1 Introduction
Several quantum many-body approaches have been developed as alternatives to DFT for electronic structure calculations, particularly in systems where strong correlation or high accuracy is essential. Methods such as GW [27, 28], dynamical mean-field theory (DMFT) [29, 30], coupled cluster theory [31, 32], and auxiliary-field quantum Monte Carlo (AFQMC) [33, 34, 35, 36] have been successfully applied to both molecular and condensed-phase systems. More recently, full configuration interaction quantum Monte Carlo (FCIQMC) [37, 38] and neural network-based quantum Monte Carlo [39, 40, 41] methods have also gained attention. However, among the quantum Monte Carlo methods, real-space fixed-node diffusion Monte Carlo (FN-DMC) remains the most widely used approach in materials science and quantum chemistry, offering a compelling balance between accuracy, scalability, and methodological maturity. Its use of explicitly correlated many-body wave functions and its favorable scaling with system size make it particularly attractive for benchmarking and systematic studies. For these reasons, this work focuses on FN-DMC and its reproducibility across a variety of independently developed community codes.
FN-DMC is an accurate state-of-the-art computational approach for solving the Schrödinger equation for a variety of systems, including molecules, solids, and surfaces. This method obtains the electronic ground-state by performing an imaginary-time evolution from a starting trial wave function . Within the Born-Oppenheimer approximation, consists of the real space positions of all the electrons. Typically, is the product of an antisymmetric function (e.g., a Slater determinant or a sum of Slater determinants[42]) and a symmetric, positive function, called the Jastrow factor[43]. The Jastrow factor is explicitly dependent on electron-electron and electron-nucleus distances, and able to directly capture a significant fraction of the electronic correlation.
@@
There are numerous QMC codes currently used for research, many of which have been under development for over a decade. Each makes somewhat different algorithmic and implementation choices, such as the use of different Jastrow factors and methods for evaluating single-particle orbitals. In this study, we compare eleven such codes and provide details of the algorithmic and implementation choices in Section S8 of the Supplementary Information (SI). This diversity raises open questions on the reproducibility of FN-DMC. If FN-DMC is to be widely accepted as a highly accurate reference method, it is important that consistent results can be obtained from these different FN-DMC codes. With this goal in mind, the present work represents a collaborative effort among the users and developers of eleven distinct FN-DMC codes, to rigorously assess the reproducibility of FN-DMC.
@@
A key obstacle to the reproducibility of FN-DMC comes from the use of non-local pseudopotentials (NLPPs), which increase the efficiency of the method for systems with heavy atoms. [107, 108, 109] While all-electron FN-DMC calculations are possible for light atoms, the computational cost increases steeply with atomic number, scaling approximately as with between 5.5 and 6.5, depending on the method details [107, 108, 110]. As a result, pseudopotentials are essential for practical FN-DMC applications involving heavier elements. NLPPs allow one to solve the Schrödinger equation solely for the valence electrons, by substituting the full local nuclear potential with a smooth non-local potential near the nuclei. In general, NLPPs hinder reproducibility in electronic structure methods, as NLPPs constructed in different ways can lead to somewhat different predictions. NLPPs are a potential source of non-reproducibility in FN-DMC even when the same NLPPs are used, because non-local pseudopotential operators create an additional sign problem in the projector beyond the one that is always present for fermionic calculations. To avoid this sign problem, these operators must be “localized”,[111] or at least partially localized,[112] on a wave-function. A natural choice is to localize them on the trial wave-function , introducing a dependence on both the determinantal and the Jastrow components of the wave function. Since the Jastrow factor is different in the different codes and its parameters are stochastically optimized, yielding possible noise and reproducibility issues, some authors choose to localize only on the determinantal component. [113, 108, 114, 115, 116] This removes the dependence on the Jastrow factor at the cost of losing the desirable property that the treatment of the pseudopotential is exact in the limit of exact . To summarize, there are currently four localization schemes: the locality approximation (LA)[117, 111], the T-move (TM) approximation[112, 118, 119], the determinant locality approximation (DLA)[113, 108, 114, 115, 116], and the determinant T-move (DTM) approximation[116]. These four schemes (LA, TM, DLA and DTM) result in somewhat different projected wave-functions and therefore different total energies of physical systems.
As computational science matures, reproducibility and transparency are increasingly recognized as critical features of robust methodology. FN-DMC, while a powerful and widely used method, has historically lacked comprehensive cross-code validation. This work takes a step toward establishing that foundation by systematically comparing the four localization algorithms across eleven FN-DMC codes (named alphabetically): Amolqc, CASINO[46], CHAMP-EU[120], CHAMP-US[121], CMQMC, PyQMC[122], QMC=Chem[123, 124], QMCPACK[48, 125], QMeCha[126], QWalk[127], and TurboRVB[47, 128]. Different forms of Jastrow factor are necessarily tested as part of this evaluation.
We specifically consider the cases of the total energies of a methane molecule, a water molecule, and a methane-water dimer, and the corresponding interaction energy. We selected the water–methane dimer as a test case not only for its modest size—which allows tight statistical convergence—but also because it spans two different interaction regimes. It involves both weak intermolecular interactions (with a binding of only ) and intramolecular energetics, enabling a sensitive probe of algorithmic consistency across codes. In addition, it is a prototype of more complex systems such as methane clathrates, important for gas storage and transportation. We show that consensus across all eleven codes can be made when utilizing the TM, DLA and DTM approximations, particularly following careful control of the discretized time step.
@@ 2 Results and Discussion
is defined as the difference between the energy of the complex, , minus the sum of the energies of the isolated water and methane monomers (see the Methods section for details on the geometries and the DMC simulation set-up). All the interaction energies are extrapolated to the zero time-step limit according to the procedure described in the SI and in Ref. [129].
@@ 3 Methods
The interaction energy of the methane-water dimer is computed by subtracting the isolated molecule energies from the methane-water complex, as defined in Eq. 1. The geometry of the dimer (shown in the SI) was obtained from Ref. [135]. The geometries of the monomers are the same as in the dimer. In this study, in order to try to achieve consistent results, all eleven codes were required to use the same correlation consistent effective core potential (ccECPs)[136, 137] and the corresponding triple-zeta basis set (ccECP-ccpVTZ), as well as a Slater-Jastrow wave function with a single Slater determinant whose orbitals are obtained from DFT calculations using the Perdew-Zunger parametrization[138] of the local-density approximation. For the methane-water dimer, this was sufficient to obtain accurate results. Some of the codes exchanged wave function data via the TREXIO library. [139] This choice ensures that any observed variation is due to implementation-level or algorithmic factors rather than differences in the choice of geometry, pseudopotential, basis set, or single-particle orbitals. Every code implements a slightly different parametrization of the Jastrow factor, but all codes include in the Jastrow factor an electron-electron (e-e), an electron-nucleus (e-n), and an electron-electron-nucleus (e-e-n) term. The variational parameters of the Jastrow factor have been optimized by minimizing either the variational energy or the variance, according to the recommended scheme within each code. The time steps employed in each simulation are in the range 0.001 to 0.1 a.u. The final estimates reported in Fig. 1 were extrapolated to the limit using the procedure described in the SI. Further details specific to each code, the schemes used to deal with the localization error, the time step extrapolation, and the tests on the size consistency error are reported in the SI.
@@ 4 Summary and Conclusions
In particular, we conducted a thorough analysis of two key obstacles to FN-DMC reproducibility, namely the use of NLPPs and finite time-step bias. We systematically compared four localization schemes, LA, TM, DLA, and DTM, for the interaction energy of the methane-water dimer and the total energies of the methane and water molecules and of the dimer. We found that agreement in the interaction energy across all eleven codes is achieved in the limit of zero time step when employing the TM, DLA, and DTM approximations. Notably, we achieve agreement within a standard deviation of meV on the interaction energy of the methane-water complex, approximately two hundred thousand times smaller than the total energy of the dimer. Larger discrepancies are observed with the LA scheme. Agreement in total energies across codes is also achieved, at sub-millihartree precision. In particular, the total energies with the TM, DLA, and DTM schemes have a standard deviation among the codes which is smaller than meV. This agreement further reinforces the reproducibility of FN-DMC.
Looking ahead, extending this cross-code effort to periodic solids would be a natural next step. However, such systems introduce additional layers of complexity—including basis set periodization, Brillouin zone sampling, and finite-size corrections—that go beyond the scope of this initial benchmark. Moreover, as only a subset of the participating codes currently support periodic boundary conditions, we deliberately focused here on molecular systems in open boundary conditions to establish a controlled but challenging comparison for FN-DMC reproducibility.
Supplementary Material
See the supplementary material for comprehensive details on the computational setup used in this study, including the geometry of the systems, descriptions of the trial wave functions, and specific parameters for each of the 11 FN-DMC codes. Additional data are provided on time-step convergence studies, localization error analysis, interaction and total energy comparisons. The file also includes technical implementation notes from each code, information on Jastrow factor optimization, and complete tables of all raw FN-DMC energies and statistical uncertainties used to generate the figures in the main text.
@@ Acknowledgements
M.B. acknowledges the computational resources from the HPC facilities of the University of Luxembourg [140] (see hpc.uni.lu). M.D. acknowledges financial support from the European Union under the LERCO project number CZ.10.03.01/00/22_003/0000003, via the Operational Programme Just Transition, and computational resources from the IT4Innovations National Supercomputing Center (e-INFRA CZ, ID: 90140). M.C. acknowledges access to French computational resources at the CEA-TGCC center through the GENCI allocation number A0150906493. J.T.K., P.R.C.K., Y.L. and L.M. were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. Part of the work of L.M. has been supported also by U.S. National Science Foundation grant DMR-2316007 and employed resources at NERSC at early stages for this project. M.C., E.S., R.S., and C.F. acknowledge partial support by the European Centre of Excellence in Exascale Computing TREX — Targeting Real Chemical Accuracy at the Exascale. This project has received funding in part from the European Union’s Horizon 2020 — Research and Innovation program — under grant agreement no. 952165. E.S., R.S., and C.F performed the calculations on the Dutch national supercomputer Snellius with the support of SURF Cooperative. K.N. acknowledges financial support from the JSPS Overseas Research Fellowships and from MEXT Leading Initiative for Excellent Young Researchers (Grant No. JPMXS0320220025) and computational resources from the Numerical Materials Simulator at National Institute for Materials Science (NIMS). L.K.W and W.W. were supported by U.S. National Science Foundation via Award No. 1931258. Y.S.A., A.Z. and D.A. acknowledge support from the European Union under the Next generation EU (projects 20222FXZ33 and P2022MC742) and from Leverhulme grant no. RPG-2020-038. A.M. and B.X.S. acknowledge support from the European Union under the “n-AQUA” European Research Council project (Grant No. 101071937). The portion of the work done by T.A.A. and C.J.U. received initial support under AFOSR (Grant No. FA9550-18-1-0095) and was completed under the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration."
@@
This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Calculations were also performed using the Cambridge Service for Data Driven Discovery (CSD3) operated by the University of Cambridge Research Computing Service (www.csd3.cam.ac.uk), provided by Dell EMC and Intel using Tier-2 funding from the Engineering and Physical Sciences Research Council (capital grant EP/T022159/1 and EP/P020259/1), and DiRAC funding from the Science and Technology Facilities Council (www.dirac.ac.uk). This work also used the ARCHER UK National Supercomputing Service (https://www.archer2.ac.uk), the United Kingdom Car Parrinello (UKCP) consortium (EP/ F036884/1). FDP, BXS, and AM acknowledge EuroHPC Joint Undertaking for awarding the project ID EHPC-REG-2024R02-130 access to Leonardo at CINECA, Italy.